# From News Sharers to Post Viewers: How Topic Diversity and Conspiracy Theories Shape Engagement With Misinformation During a Health Crisis?

### v2

@@ 2.3. Engagement With COVID-19 Misinformation on Social Media
The engagement with misinformation during health crises, particularly COVID-19, posed significant threats to the pandemic control and public safety. Several studies have scrutinized the factors that can influence engagement with COVID-19 misinformation on social media. Recognizing emotions as a significant driver of online information diffusion (Chuai and Zhao, 2022; Pröllochs et al., 2021b, a), research in this body has predominantly focused on the role of emotions during COVID-19 (Solovev and Pröllochs, 2022; Chuai et al., 2022; Freiling et al., 2021; Charquero-Ballester et al., 2021). For example, Solovev and Pröllochs (2022) conducted a large-scale computational analysis on the role of moral emotions in the spread of COVID-19 misinformation. They found that false posts with higher other-condemning emotions (e. g., anger) are more viral (i. e., more reposts) than true posts. Based on experimental design, Freiling et al. (2021) investigated the role of anxiety in misinformation belief and sharing during COVID-19, revealing that anxiety is a driving factor in believing and sharing both true and false claims. Wang and Rzeszotarski (2023) put the focus on the images and found that online posts containing COVID-19 misinformation images receive similar interactions with posts containing random images. However, COVID-19 misinformation images are shared for longer periods than non-misinformation ones. Additionally, COVID-19 misinformation, mimicking the format and language features of news and scientific reports, is likely to receive more likes (Ngai et al., 2022). For the concern that social bots may share a large amount of misinformation during COVID-19, Teng et al. (2022) analyzed different types of users in the engagement with misinformation. They found that the role of social bots in misinformation sharing is limited, and individuals with human-like behavior play a more prominent role.
Knowing the potential role of topics in engagement with misinformation, Ngai et al. (2022) focused on the specific topics in the engagement with COVID-19 vaccine misinformation. They found that safety concern was the most significant content theme but negatively associated with likes and shares. However, during COVID-19, various aspects of society and individuals became interconnected. COVID-19 misinformation covered a broad spectrum of topics, including public health, politics, medicine, vaccines, and so forth (Wang et al., 2022; Chandrasekaran et al., 2020). The multi-topic misinformation about COVID-19 carries high uncertainty and easily causes confusion to individuals, thus posing threats to people and society (Wang et al., 2022). It has been revealed that individuals can adaptively link multiple related topics and make novel connections through the mechanism of associative inference, which increases the susceptibility to misinformation (Lee et al., 2023).
@@ 2.4. Multi-Topic Context and Conspiracy Theories During COVID-19
The discussion on COVID-19 during the pandemic has not only focused on the direct health impacts but has also spilled over into various social, economic, and political domains. Within the social domain, at the beginning of the pandemic, rumors and conspiracy theories about the origin of the virus in China were rampant on social media, which exploits anti-Asian racism and hate speech on social media (Huang et al., 2023). Using COVID-19 as a weapon, malicious misinformation fosters online hate to spread quickly beyond the control of social media platforms (Velásquez et al., 2021). The rising anti-Asian racism during the COVID-19 pandemic knocked Asian businesses (Huang et al., 2023; Luca et al., 2024; Tang et al., 2023). In the economic domain, the pandemic hit the global economy, particularly impacting sectors such as travel and leisure (Wang et al., 2023b). Additionally, the scientific uncertainty surrounding the COVID-19 pandemic fuels the weaponization of COVID-19 for politicization (Kreps and Kriner, 2020). Some individuals and groups have sought to exploit the pandemic to spread extremist ideologies and radicalize others (Davies et al., 2023). The high degree of politicization within COVID-19 information contributes to polarization in attitudes towards the pandemic and even to political ideologies (Hart et al., 2020). For example, the discussion on vaccination become gradually polarized, expanding far beyond the medical discourse and entering the domain of organizational politics (Dehghan and Nagappa, 2022). Taken together, the COVID-19 pandemic is incorporated into many other societal domains to pursue specific interests, yielding significant spillover effects on the stability of societies. In this paper, we use topic diversity to denote the extent to which the information involves multiple topics and explore its impact on the engagement with misinformation on social media during COVID-19.
@@ 3.1. Dataset: MM-COVID
The data for our analysis is sourced from a publicly available repository on GitHub (Li et al., 2020). This repository provides an extensive COVID-19 dataset that consists of source news claims and their corresponding engagement on X. Initially, the repository gathers these news claims and their veracity verdicts from reputable third-party fact-checking organizations worldwide, including Snopes and Poynter, as well as official health websites. To keep the quantity of the source claims in each language, the repository filters six languages: English, Spanish, Portuguese, Hindi, French, and Italian, which enables a broad and diverse examination of misinformation across linguistic and cultural contexts. Subsequently, online posts that explicitly reference these news sources are collected from X through its advanced search API. Notably, the news sources in the dataset are linked to the original news URLs. However, due to the diverse range of news items collected from various languages and web pages, precisely parsing the original news pages and obtaining their content is challenging. Instead, the dataset provides news claims, representing the main opinions and topics.
Data preprocessing: We begin by translating all the non-English texts in source news claims and their associated posts into English using the Googletrans Python package. Subsequently, we perform text preprocessing by removing non-alphabetic and non-numerical characters, stop words, and URLs. We lemmatize all the words using the English transformer pipeline on Spacy and only consider news claims and posts that contain at least one lemmatized word. Additionally, we exclude posts that were posted before January 1, 2020. Due to the absence of release dates for source news items in the dataset, we use the first post that links to the specific source news item as its release date. In the end, we obtain news claims and associated posts. Specifically, posts are linked to false news claims, and posts are linked to true news claims.
@@ 3.2. COVID-19 Topic Spectrum
A previous study analyzed over 13 million English COVID-19 posts and identified 26 topics within 10 broad themes, including source, prevention, spread and growth, treatment and recovery, impact on the economy and markets, impact on health care sector, and government response, political impact, and racism (Chandrasekaran et al., 2020). This research also provided keywords corresponding to each topic. To ensure a comprehensive understanding of the topic spectrum surrounding COVID-19 discussions, we utilize these keywords as seeds to identify similar words and construct a topic lexicon based on a Word2Vec model. Notably, after reexamination, we remove the topic about locations that have no specific meanings and consider the remaining 25 topics across 10 themes (Table 1).
@@
Impact on health care sector | Impact on hospitals and clinics, policy changes, |
@@
Political impact | Political impact |
@@
Word2Vec model: To train the Word2Vec model, we additionally collect a large-scale COVID-19 dataset from a public repository on Kaggle.111https://www.kaggle.com/datasets/smid80/coronavirus-covid19-tweets This dataset contains posts in multiple languages, posted from March 9 to April 30, 2020 on X. Given the size of the dataset, we only filter English posts and do not involve the translation process. By combining this dataset with MM-COVID, we obtain a total of English posts. After removing non-alphabetic and non-numerical characters as well as URLs in the posts, we utilize the refined corpus to train a Word2vec model with a vector size of 200 using the gensim library.222The model is trained on the High Performance Computing platform at the University of Luxembourg. The examples of the keywords and their ten most similar words are shown in Table 2.
@@
Topic lexicon: We expand our set of topic keywords by incorporating words that exhibit similarities of more than 0.6 with the initial topic seeds. This process yields 828 additional topic keywords. Next, we conduct a manual review of these keywords, identifying that 90.6% of them are indeed relevant to the corresponding topic seeds and effectively represent the associated topics. Consequently, we integrate the topic seeds with these related keywords to create a comprehensive topic lexicon.
@@ 3.3. Topic Diversity
Topic diversity indicates the extent to which the content involves multiple topics. To give a sense of topic diversity, we construct topic co-occurrence networks for true and false news claims. Specifically, we calculate the frequency of topic co-occurrence of any two topics out of 25 topics in true and false news claims, and retain the co-occurrence edges with frequencies exceeding the median threshold. In the network of true news claims (Fig. 1(a)), the number of edges is 107, while in the network of false news claims (Fig. 1(b)), the number of edges is 167. Compared to true news claims, false news claims contain 56.1% more co-occurrence edges and are more interconnected, indicating higher interconnectedness and suggesting greater topic diversity in false news claims. Building upon previous research that used the Gini coefficient to measure the inequality of topic distribution in the content (Miani et al., 2022), we quantitatively measure topic diversity () as:
@@ 3.4. COVID-19 Conspiracy Theories
Conspiracy theories have been often thought of as a subset of misinformation (Hanley et al., 2023). Following the method commonly adopted by previous related research (Erokhin et al., 2022; Moffitt et al., 2021; Sharma et al., 2022), we select a set of keywords (phrases) from the topic lexicon to identify the conspiracy theories embedded in the false news items and associated posts. Specifically, we consider eight COVID-19 conspiracy theory categories that connect to at least one topic in our analysis: the sources and origins of the virus that caused the outbreak of COVID-19 (VirusOrigin), the connection between 5G and COVID-19 (5G), the role of big pharmaceutical companies (BigPharma) and Bill Gates (BillGates) in the spread of COVID-19, the scope of the pandemic is exaggerated (Exaggeration), the claim that hospitals were actually empty (FilmYourHospital), the idea that the genetically modified crops led to the emergence of COVID-19 (GMO), and skepticism regarding the ineffectiveness and side-effects of vaccines (Vaccines).
As shown in Table 3, we match the keywords with the lemmatized words in each news claim or post, and for phrases, we directly match them with the original translated text. We find that false news claims and false news posts incorporate keywords (phrases) indicating conspiracy theories. Given that the share of conspiracy posts ranges from 0.6% to 18% (Erokhin et al., 2022), the 6% share of posts containing conspiracy theories in our dataset is reasonable. Upon a manual check of news items that contain keywords or phrases related to conspiracy theories, we find that the accuracy is 92.1%, which suggests that the selected keywords (phrases) can effectively identify the conspiracy theories integrated into the misinformation (Erokhin et al., 2022; Moffitt et al., 2021).
@@
3.5. The Engagement From Source News to Posts on X
On the platform of X, users can interact with posts they view through reposting, liking, or replying. Existing research often uses the number of reposts to study engagement with misinformation on social media (Vosoughi et al., 2018; Chuai and Zhao, 2022; Pröllochs et al., 2021b; Drolsbach and Pröllochs, 2023; Chuai et al., 2024). Additionally, the number of likes and the number of replies are also important indicators of engagement with online posts (Chuai et al., 2024; Papakyriakopoulos and Goodman, 2022). Notably, news items outside of X enter the social environment through news sharers on X. Multiple posts on X can link to the same original news items, which increases the visibility of the news items. However, few studies examine engagement with misinformation from the perspective of source news items to posts. In this paper, we consider user engagement at two levels: news sharer level and post viewer level (Fig. 2). At news sharer level, news items are shared via posts on X. The engagement metrics include the number of posts and the lifetime of news items from the creation of the first post to the creation of the latest post. At post viewer level, engagement is measured by how viewers interact with these posts. Post viewers can repost, like, and reply to the posts that are created by news viewers and contain links to the original news items.
@@ 3.6. Empirical Models
We employ multiple explanatory regression models to explore the research questions separately. To this end, we define the following dependent variables to measure engagement:
@@
The key independent variables include:
@@
We also control for the variables from post and user characteristics:
: A continuous-like count variable indicating the number of words in the text of the news claim or post.
: A dummy variable indicating whether the post has media elements ( 1) or not ( 0).
@@
Additionally, emotions are considered as a main driver of online information diffusion (Chuai and Zhao, 2022; Pröllochs et al., 2021b, a). Our analysis considers six cross-cultural emotions – anger, disgust, fear, joy, sadness, and surprise – out of the eight basic emotions (Chuai and Zhao, 2022; Plutchik, 1991). We employ the commonly used NRC emotion lexicon to count emotion words in news claims and posts. Notably, 92.1% news claims and 95% posts contain at least one emotion word. Subsequently, we calculate the weight of each emotion in the specific text content based on the number of corresponding emotion words. If the news claims (posts) have no emotion words, all the weights of emotions are set to 0. The descriptive statistics for the variables are shown in Table 4.
@@
Topic diversity (RQ1 & RQ2): We use as the dependent variable to analyze the topic diversity in true and false news from original news claims to online posts. In terms of the claims, the linear regression model that incorporates the main independent variables of and is specified as:
@@
where indicates news-specific random effects. Additionally, we also incorporate , poster characteristics as control variables in the model.
Engagement at news sharer level – lifetime and post count (RQ3 & RQ4): The first layer of engagement is from source news claims to news posts on X. We examine how topic diversity and conspiracy theories influence engagement with misinformation in terms of news lifetime and post count. First, we take as the dependent variable and specify a linear regression model as follows:
@@
Engagement at post viewer level – reposts, likes, and replies (RQ3 & RQ4): To examine the effects of topic diversity and conspiracy theories on social engagement (i. e., reposts, likes, and replies) with misinformation on X/Twitter, we specify a negative binomial regression model with language-specific fixed effects, month-year fixed effects, and news-specific random effects:
where indicates all the independent variables, and signifies the dependent variables, including , and . All the continuous variables in the above models are z-standardized.

### v3

@@ 2.3. Engagement With COVID-19 Misinformation on Social Media
The engagement with misinformation during health crises, particularly COVID-19, poses significant threats to crisis control and public safety. Several studies have scrutinized the factors that can influence engagement with COVID-19 misinformation on social media. Recognizing emotions as a significant driver of online information diffusion (Chuai and Zhao, 2022; Pröllochs et al., 2021b, a), research in this body has predominantly focused on the role of emotions during COVID-19 (Solovev and Pröllochs, 2022; Chuai et al., 2022; Freiling et al., 2023; Charquero-Ballester et al., 2021). For example, Solovev and Pröllochs (2022) conducted a large-scale computational analysis on the role of moral emotions in the spread of COVID-19 misinformation. They found that false posts with higher other-condemning emotions (e. g., anger) are more viral (i. e., more reposts) than true posts. Based on experimental design, Freiling et al. (2023) investigated the role of anxiety in misinformation belief and sharing during COVID-19, revealing that anxiety is a driving factor in believing and sharing both true and false claims. Wang and Rzeszotarski (2023) put the focus on the images and found that online posts containing COVID-19 misinformation images receive similar interactions with posts containing random images. However, COVID-19 misinformation images are shared for longer periods than non-misinformation ones. Additionally, COVID-19 misinformation, mimicking the format and language features of news and scientific reports, is likely to receive more likes (Ngai et al., 2022). For the concern that social bots may share a large amount of misinformation during COVID-19, Teng et al. (2022) analyzed different types of users in the engagement with misinformation. They found that the role of social bots in misinformation sharing is limited, and individuals with human-like behavior play a more prominent role.
Knowing the potential role of topics in engagement with misinformation, Ngai et al. (2022) focused on the specific topics in the engagement with COVID-19 vaccine misinformation. They found that safety concerns are the most significant content theme but are negatively associated with likes and shares. However, during COVID-19, various aspects of society became interconnected. COVID-19 misinformation covered a broad spectrum of topics, including public health, politics, medicine, vaccines, and so forth (Wang et al., 2022; Chandrasekaran et al., 2020). The multi-topic misinformation about COVID-19 carries high uncertainty and easily causes confusion to individuals, thus posing threats to people and society (Wang et al., 2022). It has been revealed that individuals can adaptively link multiple related topics and make novel connections through the mechanism of associative inference, which increases the susceptibility to misinformation (Lee et al., 2023).
@@ 2.4. Multi-Topic Context and Conspiracy Theories During COVID-19
The discussion on COVID-19 during the pandemic not only focused on the direct health impacts but also spilled over into various social, economic, and political domains. Within the social domain, at the beginning of the pandemic, rumors and conspiracy theories about the origin of the virus in China were rampant on social media, exploiting anti-Asian racism and hate speech on social media (Huang et al., 2023). Using COVID-19 as a weapon, malicious misinformation fostered online hate to spread quickly beyond the control of social media platforms (Velásquez et al., 2021). The rising anti-Asian racism during the COVID-19 pandemic knocked Asian businesses (Huang et al., 2023; Luca et al., 2024; Tang et al., 2024). In the economic domain, the pandemic hit the global economy, particularly impacting sectors such as travel and leisure (Wang et al., 2023). Additionally, the scientific uncertainty surrounding the COVID-19 pandemic fueled the weaponization of COVID-19 for politicization (Kreps and Kriner, 2020). Some individuals and groups have sought to exploit the pandemic to spread extremist ideologies and radicalize others (Davies et al., 2023). The high degree of politicization within COVID-19 information contributed to polarization in attitudes towards the pandemic and even to political ideologies (Hart et al., 2020). For instance, the debate over vaccination became increasingly polarized during the pandemic, expanding far beyond the medical discourse and entering the domain of organizational politics (Dehghan and Nagappa, 2022). Taken together, the COVID-19 pandemic has been incorporated into many other societal domains to pursue specific interests, yielding significant spillover effects on the stability of societies. In this paper, we use topic diversity to denote the extent to which the information involves multiple topics and explore its impact on the engagement with misinformation on social media during COVID-19.
@@ 3.1. Dataset: MM-COVID
The data for our analysis is sourced from a publicly available repository on GitHub (Li et al., 2020). This repository provides an extensive COVID-19 dataset that consists of source news claims and their corresponding engagement on X. Initially, the repository gathers these news claims and their veracity verdicts from popular third-party fact-checking organizations worldwide, including Snopes and Poynter, as well as official health websites. To keep the quantity of the source claims in each language, the repository filters six languages: English, Spanish, Portuguese, Hindi, French, and Italian. This enables a broad and diverse examination of misinformation across linguistic and cultural contexts. Subsequently, online posts that explicitly reference these news sources are collected from X through its advanced search API. Notably, while the news sources in the dataset are linked to the original news URLs, precisely parsing the original news pages and obtaining their content is challenging due to the diverse range of news items collected from various languages and web pages. Instead, the dataset provides news claims, representing the main opinions and topics.
Data preprocessing: We begin by translating all the non-English texts in source news claims and their associated posts into English using Google Translate API via Googletrans.111Googletans library is available at https://github.com/ssut/py-googletrans. It is a free and unlimited Python library that implements Google Translate API and uses the same servers that translate.google.com uses. Google Translate has a high performance in translating between English and non-English languages, retaining overall meaning in 82.5% of the translations (Taira et al., 2021; Sebo and de Lucia, 2024). Subsequently, we perform text preprocessing by removing non-alphabetic and non-numerical characters, stop words, and URLs. We lemmatize all the words using the English transformer pipeline on Spacy and only consider news claims and posts that contain at least one lemmatized word. Additionally, we exclude posts that were posted before January 1, 2020. Due to the absence of release dates for source news items in the dataset, we use the first post that links to the specific source news item as its release date. In the end, we obtain news claims and associated posts. Specifically, posts are linked to false news claims, and posts are linked to true news claims.
@@ 3.2. COVID-19 Topic Spectrum
A previous study analyzed over 13 million English COVID-19 posts and identified 26 topics within 10 broad themes, including source, prevention, spread and growth, treatment and recovery, impact on the economy and markets, impact on health care sector, and government response, political impact, and racism (Chandrasekaran et al., 2020). This research also provided keywords corresponding to each topic. To ensure a comprehensive understanding of the topic spectrum surrounding COVID-19 discussions, we utilize these keywords as seeds to identify similar words and construct a topic lexicon based on a Word2Vec model. Upon reevaluating the associations between topic categories and their themes, we find that the keywords in the topic “Hotspots and location” relate solely to places, such as “cities” and “location,” not adequately reflecting the theme “Spread and growth” in terms of transmission modes or case numbers. Therefore, in the main analysis, we remove the topic “Hotspots and location” and focus on the remaining 25 topics across 10 themes (Table 1).222In a subsequent robustness check, we conduct the analysis again using all the original topics and keywords. The results are robust and consistently support our findings. The full estimation results are reported in Suppl. S3.
@@
| Impact on health care sector | Impact on hospitals and clinics, health policy, |
@@
| Political impacts | Political impacts |
@@
Word2Vec model: To train the Word2Vec model, we additionally collect a large-scale COVID-19 dataset from a public repository on Kaggle.333The dataset is available at https://www.kaggle.com/datasets/smid80/coronavirus-covid19-tweets. This dataset contains posts in multiple languages, posted from March 9 to April 30, 2020 on X. Given the size of the dataset, we only filter English posts and do not involve the translation process. By combining this dataset with MM-COVID, we obtain a total of English posts. After removing non-alphabetic and non-numerical characters as well as URLs in the posts, we utilize the refined corpus to train a Word2vec model with a vector size of 200 using the gensim library.444The model is trained on the High Performance Computing platform at the University of Luxembourg. The examples of the keywords and their ten most similar words are shown in Table 2.
Topic lexicon: We expand our set of topic keywords by incorporating words that exhibit cosine similarities of more than 0.6 with at least one of the initial topic seeds. This process yields 828 additional topic keywords, and they are automatically categorized into the same topics as their corresponding topic seeds. Additionally, we conduct a manual review of these keywords and remove ambiguous terms and abbreviations. For example, we remove the keyword “reason” from the topic “Alternative causes,” as it does not specify a particular alternative cause for COVID-19, whereas the keyword “chinavirus” is an effective term for this topic category. Similarly, we remove ambiguous abbreviations such as “pres” in the topic “Political impacts.” Ultimately, we retain 90.6% of the 828 keywords and integrate them with the topic seeds to create an extensive topic lexicon.555We repeat our analysis using the keywords without manual validation. The results remain consistent, further enhancing the robustness of our findings. The full estimation results are reported in Suppl. S3.
@@ 3.3. Topic Diversity
Topic diversity indicates the extent to which the content involves multiple topics. To give a sense of topic diversity, we construct topic co-occurrence networks for true and false news claims. Specifically, we calculate the frequency of topic co-occurrence of any two topics out of 25 topics in true and false news claims, and retain the co-occurrence edges with frequencies exceeding the median threshold. In the network of true news claims (Fig. 1(a)), the number of edges is 107, while in the network of false news claims (Fig. 1(b)), the number of edges is 167. Compared to true news claims, false news claims contain 56.1% more co-occurrence edges and are more interconnected, indicating higher interconnectedness and suggesting greater topic diversity in false news claims. For example, the topic “Political impacts” is connected to 5 topics in true news claims, whereas it connects to 10 additional topics, such as “Death reports” and “Drugs and vaccines,” in false news claims (Fig. 1(c)).
Building upon previous research that used the Gini coefficient to measure the inequality of topic distribution in the content (Miani et al., 2022), we quantitatively measure topic diversity () as:
@@ 3.4. COVID-19 Conspiracy Theories
The outbreak of COVID-19 has fueled widespread conspiracy theories with misleading claims about its origin, treatment, and other related topics. We summarize the conspiracy theories identified in previous research (Erokhin et al., 2022; Moffitt et al., 2021; Sharma et al., 2022b) and use a set of keywords and phrases to identify conspiracy theories embedded in false news claims and subsequent posts. Specifically, we consider eight COVID-19 conspiracy theory categories that connect to at least one topic in our analysis: the sources and origins of the virus that caused the outbreak of COVID-19 (VirusOrigin), the connection between 5G and COVID-19 (5G), the role of big pharmaceutical companies (BigPharma) and Bill Gates (BillGates) in the spread of COVID-19, the claim that the scope of the pandemic was exaggerated (Exaggeration), the claim that hospitals were actually empty (FilmYourHospital), the idea that the genetically modified crops led to the emergence of COVID-19 (GMO), and the skepticism regarding the ineffectiveness and side-effects of vaccines (Vaccines).
We combine the conspiracy-specific terms adopted by prior studies (Erokhin et al., 2022; Moffitt et al., 2021) and compile an extensive list of keywords and phrases to identify the eight conspiracy theories (Table 3). Subsequently, we match these keywords with the lemmatized words in each news claim or post, and for phrases, we directly match them with the original translated text. We find that false news claims and false news posts incorporate keywords (phrases) indicating conspiracy theories. Given that the shares of conspiracy posts range from 0.6% to 18% in previous studies (Erokhin et al., 2022), the 6% share of posts containing conspiracy theories in our dataset is reasonable. Upon a manual check of the source news items that contain keywords or phrases related to conspiracy theories, we find that the accuracy is 92.1%. Therefore, the selected keywords (phrases) can effectively identify conspiracy theories in our dataset.
@@
3.5. User Engagement on X
On the platform of X, users can interact with posts they view through reposting, liking, or replying. Existing research typically uses the number of reposts to study engagement with misinformation on social media (Vosoughi et al., 2018; Chuai and Zhao, 2022; Pröllochs et al., 2021b; Drolsbach and Pröllochs, 2023; Chuai et al., 2024b). Additionally, the number of likes and the number of replies are also important indicators of engagement with online posts (Chuai et al., 2024b; Papakyriakopoulos and Goodman, 2022). However, few studies examine engagement with misinformation from the perspective of news source sharing. Multiple posts on X can link to the same source news items, which increases the visibility of the news items. In this paper, we consider user engagement at two levels: news source sharer level and post viewer level (Fig. 2). At news source sharer level, news items are shared by news source sharers via posts on X. The engagement metrics include the number of posts and the lifetime of news items from the creation of the first post to the creation of the latest post. At post viewer level, engagement is measured by how viewers interact with these posts. Post viewers can repost, like, and reply to posts that are created by news source sharers and contain links to the original news items.
@@ 3.6. Empirical Models
We employ multiple explanatory regression models to explore our research questions separately. To this end, we define the following variables to measure engagement at news source sharer level and post viewer level.
The engagement at news source sharer level:
@@
Additionally, we define the following key variables to address our research questions:
@@
We also control for the variables from content characteristics that can affect user engagement:
: A continuous-like count variable indicating the number of words in the text of the news claim or post. This control variable is commonly used to account for the content richness (Miani et al., 2022).
: A dummy variable indicating whether the post has media elements ( 1) or not ( 0). Previous research has found that media elements, such as images and videos, can facilitate user engagement (Zhou et al., 2021).
Emotions are considered to be the main driver of online information diffusion (Chuai and Zhao, 2022; Pröllochs et al., 2021b, a). Our analysis considers six cross-cultural emotions – anger, disgust, fear, joy, sadness, and surprise – out of the eight basic emotions (Chuai and Zhao, 2022; Plutchik, 1991). We employ the commonly used NRC emotion lexicon to count emotion words in news claims and posts. Notably, 92.1% news claims and 95% posts contain at least one emotion word. Subsequently, we calculate the weight of each emotion in the specific text content based on the number of corresponding emotion words. If the news claims (posts) have no emotion words, all the weights of emotions are set to 0.
Finally, the posters on social media play an important role in driving engagement with their posts (Vosoughi et al., 2018; Chuai and Zhao, 2022). We use the following four variables to account for poster characteristics:
@@
The descriptive statistics for the above variables are shown in Table 4.
@@
Topic diversity (RQ1.1 & RQ2.1): We use as the dependent variable to analyze topic diversity in true and false news from original news claims to online posts. In terms of claims, the linear regression model that incorporates the main independent variables of and is specified as:
@@
where indicates news-specific random effects. Additionally, we incorporate , poster characteristics as control variables. In the above two models, the coefficient estimate of () is to address RQ1.1, and the coefficient estimate of () is to address RQ2.1.
Engagement at news source sharer level – news lifetime and post count (RQ1.2 & RQ2.2): The first layer of engagement is from source news claims to news posts on X. We examine how topic diversity and conspiracy theories influence engagement with misinformation in terms of news lifetime and post count. First, we take as the dependent variable and specify a linear regression model as follows:
@@
Engagement at post viewer level – reposts, likes, and replies (RQ1.2 & RQ2.2): To examine the effects of topic diversity and conspiracy theories on social engagement (i. e., reposts, likes, and replies) with misinformation on X, we specify a negative binomial regression model with language-specific fixed effects, month-year fixed effects, and news-specific random effects:
where indicates all the independent variables, and signifies the dependent variables, including , and . In the above models for engagement, the coefficient estimates of (), (), and () are to address RQ1.2, and the coefficient estimate of () is to address RQ2.2. All the continuous variables in these models are z-standardized.